# A Survey on Heterogeneous Computing Using SmartNICs and Emerging Data Processing Units (Expanded Preprint)

### v1

@@ Abstract
The emergence of new, off-path smart network cards (SmartNICs), known generally as Data Processing Units (DPU), has opened a wide range of research opportunities. Of particular interest is the use of these and related devices in tandem with their host’s CPU, creating a heterogeneous computing system with new properties and strengths to be explored, capable of accelerating a wide variety of workloads. This survey begins by providing background information to this new field, such as discussing its origins, its motivations and challenges, listing a few of the current market offerings for DPUs, and providing some brief information about the major programming languages and frameworks for using them. Then, we review and categorize a number of recent works in the field, covering a wide variety of studies, benchmarks, and application areas such as in data center infrastructure, commercial uses, and AI and ML acceleration.
@@ 1 Introduction
The computational hardware landscape is changing rapidly with the advent of new processors and networking equipment. In this survey, we focus on academic and industrial research on heterogeneous computing using an emerging class of network computing devices known as data processing units (DPU), in a way that is more accessible to those less familiar with networking devices, in contrast with expert-oriented surveys such as the recent work of Kfoury et al. [49]. As network interface cards (NIC) led to the development of smart network interface cards (SmartNIC), we see similar advancements in the field of SmartNICs. Recent developments in SmartNICs by companies like Nvidia, Amazon, Microsoft (fungible acquisition) and AMD (Pensando acquisition) have led to the new “DPU” nomenclature.
A distinguishing aspect of a DPU is that it integrates a general-purpose CPU, capable of running an operating system, with DRAM and storage, into the high-speed NIC datapath of a SmartNIC. Thus, modern DPUs have a higher data throughput of 100 Gbps/400 Gbps compared to traditional SmartNICs, although the CPU cores in a DPU are generally less powerful than the host’s CPU cores. To provide an example, NVidia’s Mellanox Bluefield-2 is considered a DPU.
@@
Note that while we use the term ‘DPU’ to refer to a class of advanced SmartNICs, as described, these are also known by other names, such as “Distributed Processing Unit” and “Infrastructure Processing Unit” (not to be confused with an “Intelligence Processing Unit”, which is a fundamentally different device); nomenclature is inconsistent, and depends upon the vendor and product specialization. Furthermore, the term ‘Data Processing Unit’ itself has become overloaded, as it has also been used to describe other devices, such as Processing-in-Memory (PIM) devices, which are not discussed in this work, and Deep-Learning Processing Units, which are not SmartNICs.
DPU SmartNICs are typically used in place of a server’s standard SmartNICs, which connect it to the network, and constitute an integral component of the conventional data center’s infrastructure resources, along with the CPU, memory, storage, and networking constituents. The physical hardware implementation of a DPU can consist of either System-on-chip, FPGA, or hybrid architectures.
@@
The development of advanced SmartNICs, or DPUs, was started by large cloud service providers (“hyperscalers”) such as Amazon who realized that around 30% of processor cores were being utilized for infrastructure processing [90]. This motivated cloud vendors to develop advanced SmartNICs to offload infrastructure-related tasks related to networking, security, and storage from host CPUs, so that host CPU cores can focus on users’ application workloads. According to the projection of Dell’Oro Group, the traditional NIC market share will be overtaken by SmartNICs by 2027 [79].
@@
There are many different reasons for developers and professionals to use SmartNICs, the most obvious of which is its intended use as a network infrastructure support device. For example, one such use case is to perform network offload, where networking tasks, such as packet processing, are delegated to SmartNICs by host CPUs. Their utility extends beyond basic networking functions and is also capable of encompassing security offloading, which is achieved by accelerating cryptographic operations such as those common in a Secure Sockets Layer (SSL), Transport Layer Security (TLS), and encryption. Similarly, SmartNICs are used to offload storage-related operations, as well as to support virtualization tasks.
@@
However, data center infrastructure is not the sole application domain in which SmartNICs are applied, which is the main motivation of this paper. This survey focuses on works exploring the use of DPUs for in-network computing and applications acceleration, as well as infrastructure acceleration, though we highlight the usage of the DPU as a computational and communication accelerator for non-infrastructure-related applications. Within each of these domains, such accelerations are made possible primarily by two things: 1) acceleration and offload of computations and networking functions, and 2) isolation between host and DPU. Both these and the evolution of SmartNICs will be discussed in Section 2.
Thereafter, the remainder of the survey will be organized as follows: Section 3 discusses the motivation behind the development of DPUs and the challenges encountered in leveraging this emerging class of networking hardware; Section 4 presents the variety of DPU offerings by technology vendors; Section 5 presents works regarding programming tools and frameworks for developing application software targeted to DPUs; and Section 6 presents works regarding various studies, benchmarks, and applications of DPUs, ranging from infrastructure to offloading commercial applications, including AI and ML.
@@ 2 Background
For those unfamiliar with SmartNICs, a certain amount of background information will be helpful in understanding the works presented later on.
@@ 2.1 Evolution of NICs and DPUs
Figure 1: The 2021 ServeTheHome NIC Continuum, a useful classification of existing NIC hardware, and which doubles as a representation of its evolution over time [47].
@@
The ‘Foundational NIC’ shown in Figure 1 represents the fundamental level of a network interface. Nearly all contemporary NICs possess certain rudimentary offloads, such as IPv4/IPv6 and TCP/UDP checksum offloads. Foundational NICs are specifically engineered to facilitate low-cost network ports, eschewing many of the advanced offload features that contribute to increased cost of more sophisticated devices. Foundational NICs retain their significance within the industry, but are insufficient in many cases, particularly at higher data rates where the escalation in data processing necessitates enhanced computational capabilities.
Consequently, the prevailing NICs are predominantly mid-tier ‘Offload NICs’, which generally provide networking speeds of about 100Gbps and above [47]. In order to design NICs with such speeds, it becomes necessary to offload more network support functions, generally in the form of on-board hardware accelerators (ASICs, or application-specific integrated circuits) [47] - otherwise the host CPU would be quite busy trying to keep up.
‘SmartNICs’, however, go beyond simple offloads. These are NICs that possess programmable pipelines, allowing systems and applications to tweak and configure offloads more flexibly. The primary purpose of a SmartNIC, as being part of a server device, is to reduce the burden imposed on the host CPU [47]. “Smart”, in the name, of course suggests a certain programmability; at the very least, an increased ability to handle complex tasks - both current, and being adaptable to future needs [87].
A ‘DPU’ differs from this paradigm, however, in that it can hardly be considered part of the host server - it is, in and of itself, a mini network endpoint [47, 4], providing ‘off-path’ capabilities (we will define off-path later). DPUs represent a more sophisticated iteration of SmartNICs. The concept encompasses the former functionality and extends beyond it, possessing not only ASIC offload capabilities and the flexible programmable pipelines of a SmartNIC, but also its own memory, storage, and CPUs. This design enables them to not only perform network tasks, but to participate in data processing tasks - further freeing up the host’s compute resources, all while consuming less power than the host [4]. While ServeTheHome further differentiates devices incorporating large Field-Programmable Gate Arrays (FPGA), we lump these in with DPUs for the purposes of this survey.
@@ 2.2 Computational Acceleration and Offload
DPUs are designed to address the increasing burden on CPUs in data centers caused by the massive overhead of infrastructure processing [20, 66]. A DPU, being ‘close to’ the network, of course boasts better network I/O processing performance compared to the host, making it ideal for accelerating network stack and packet processing, session management, distributed storage and file system client execution, erasure coding, network security layers, and even acting as a virtual switch [28, 58, 21]. DPUs have also been leveraged in network-edge scenarios for applications such as network awareness, power-efficient edge-to-cloud continuum, etc. [5, 63]. The goal of such offloads, of course, is to free up the host’s resources to focus on applications processing tasks and improve system-wide performance by reducing network and data center overhead [82].
And, while DPU acceleration typically refers to offloading these types of network and infrastructure tasks using on-board hardware accelerators (ASICs) [66, 63], this survey especially points out those works which instead show that one can offload generic computations and data manipulation to a DPU, including work in the applications layer - moving virtually any kind of processing from the host to the DPU’s CPUs. Some notable examples from the papers we review include big data processes, machine learning functions and inference, molecular dynamics, and various other forms of scientific computing. A simple illustration of what offloading to a DPU might look like is provided in Figure 2.
@@ 2.3 Data Path
When discussing ‘data paths’ in terms of advanced SmartNICs such as DPUs, the term may be more specific than the typical usage. In general a data path describes the steps through which data travels as it is moved, analyzed, transformed, and processed within a workflow. In a DPU, however, the data path usually traverses both the host and the DPU, and typically can be described one of two ways: ‘on-path’ or ‘off-path’, as shown in Table 1. Understanding the difference between the two is helpful in understanding the various use-cases of SmartNICs and DPUs.
On-path SmartNICs have CPU cores that are situated on the communication path of the network packets traveling to the host, and possess the ability to manipulate each incoming or outgoing packet rapidly, but too much additional processing assigned to these cores can bottleneck the communication path, because all network traffic must pass through them. Examples of on-path SmartNICs include LiquidIO II CN2350 and CN2360.
By contrast, off-path SmartNICs (typically DPUs) have cores that are not in the direct line of traffic to the host. Their internal NIC can use customizable network forwarding rules to sort between packets that should be sent to the DPU cores first for pre-processing, and those that should be forwarded directly to the host instead. However, direct control over packets and memory will be limited, and have overheads - any off-path processing which occurs lengthens the data path, and thus time, for packets to make it to the host [81]. Examples of off-path SmartNICs include the Mellanox BlueField and the Broadcom Stingray [60]. In short, a ‘DPU’ is a class of device with certain advanced capabilities, which is usually an ‘off-path’ device, a term which describes both the physical network path within the device, and a paradigm for how the data flows through it.
Many research projects take advantage the flexibility of off-path systems, analyzing the data path and data flow of a given application and experimenting with using a SmartNIC to optimize its efficiency, or save on data pre-processing steps.
@@ 2.4 Isolation
First, amid the many services, or tenants, which a server node often has to support, being able to separate, or isolate, many of the infrastructure tasks from the tenant or application workload, as well as separating those tenant workloads from each other, has the capacity to improve performance, not only because it redistributes compute resources, but also because it reduces interference between tasks; CPU context switching can have a high overhead cost, particularly due to cache pollution, and task segregation can help reduce that cost. This problem only worsens when you consider the myriad of additional ASIC resources over which there might be contention [30].
Secondly, a DPU can be used as an extra security layer. Because it functions as a separate device which can process network packets before forwarding them to the host, it provides an added layer of hardware isolation between the network and the host server, effectively isolating infrastructure and other functions performed on the SmartNIC from those performed on the host [93], which usually amounts to separating control plane operations from data plane operations, or infrastructure operations from tenant operations, or network access from user applications, thus further protecting the data center from nefarious tenants. And, with more security operations offloaded to ASICs, one could reasonably say that the more computation one does in silicon, instead of software, the more secure the system will be [89].
3 Motivation and Challenges
Approximately 30% of the work done in a data center is in handling network infrastructure tasks (colloquially called an “infrastructure tax”), meaning that it requires almost half again as many servers to get the intended work done; this severely bloats the size and cost of data centers. DPUs are increasingly seen as a solution to this problem, and CPUs as insufficient [94].
Heterogeneous computing using emerging Data Processing Units (DPUs) is driven in particular by the necessity for specialized acceleration, enhanced performance, and energy efficiency. DPUs, which are specifically designed for particular tasks, optimize the effectiveness of the system by transferring specialized workloads from the host CPU, providing a more energy-efficient solution for data center infrastructure workloads, which also improves cost-efficiency. However, since DPUs have a general-use processor that is also more energy efficient [21], they are thus capable of accommodating a wide range of workloads and enhancing the overall utilization of resources. The adaptability and flexibility of heterogeneous computing with DPUs is intended to address the evolving requirements of modern applications in this way [15].
The motivation behind the creation of these smart network devices, however, goes deeper. As the world turns to cloud technologies and flexibly hosted web services, there is an increasing need for software-defined networking (SDN) - systems capable of not only specialized, but customizable networking configurations within data centers and supercomputers [12] - perhaps for both infrastructure and tenant applications alike. DPUs will be part of the future infrastructure powering services such as augmented reality, smart cities, edge computing, big-data analytics, and more [89].
A smarter NIC, of course, comes with a higher price point, and whether or not that cost comes out to be worth it is a matter of balance, requiring the consideration of many factors, for a given data center - not least of which, the potential for millions of dollars worth of electricity and capital savings for a large data center [21]. However, the desire for software-reconfigurable optimization within the data center may be king, given the directions in which technology is advancing - so, DPU adoption is expected to keep increasing [41].

### v2

@@ Abstract
The emergence of new, off-path smart network cards (SmartNICs), known generally as Data Processing Units (DPU), has opened a wide range of research opportunities. Of particular interest is the use of these and related devices in tandem with their host’s CPU, creating a heterogeneous computing system with new properties and strengths to be explored, capable of accelerating a wide variety of workloads. This survey begins by providing the motivation and relevant background information for this new field, including its origins, a few current hardware offerings, major programming languages and frameworks for using them, and associated challenges. We then review and categorize a number of recent works in the field, covering a wide variety of studies, benchmarks, and application areas, such as data center infrastructure, commercial uses, and AI and ML acceleration. We conclude with a few observations.
keywords:
DPU , data processing units , SmartNIC , smart network cards , off-path , heterogeneous , acceleration , offload , infrastructure , parallel , distributed
journal: Future Generation Computer Systems
[…]
[MST]organization=Department of Computer Science, Missouri University of Science and Technology,addressline=500 West 15th Street, 325 Computer Science Bldg., city=Rolla, postcode=65409, state=Missouri, country=USA
@@ 1 Introduction
The computational hardware landscape is changing rapidly with the advent of new processors and networking equipment. In this survey, we review academic and industrial research using an emerging class of advanced network computing devices known as data processing units (DPU), with a particular focus on their use in heterogeneous and parallel computing applications, in contrast to more general surveys, such as the recent work of Kfoury et al. [1]. Such preexisting works are also more expert-oriented and lack the applications focus of ours; this survey is instead intended to be very accessible to those less familiar with networking devices, but who want to use them in applications.
As network interface cards (NIC) led to the development of smart network interface cards (SmartNIC), we see similar advancements in the field of SmartNICs. Recent developments in SmartNICs by companies like NVIDIA, Amazon, Microsoft (Fungible acquisition) and AMD (Pensando acquisition) have led to the new “DPU” nomenclature.
A DPU is thus essentially a class of advanced SmartNICs, with distinguishing aspects such as typically integrating a general-purpose CPU capable of running an operating system (OS), with DRAM and storage, into the high-speed NIC datapath of a SmartNIC. However, the CPU cores in a DPU are generally cheaper and less powerful than the host’s CPU cores [2], and some DPUs incorporate powerful FPGAs in place of or in addition to CPUs. DPUs also have a high data throughput, often 100-400 Gbps. To provide an example, NVIDIA’s Mellanox BlueField devices are considered to be DPUs, and their flagship DPU product is the BlueField-3; NVIDIA is one of the current industry leaders and a trend-setter in our line of research. Other examples are listed in Section 2.7. DPUs are typically used in place of a server’s standard SmartNICs, which connect it to the network; they are an increasingly integral component of conventional data centers’ infrastructure. For this survey’s heterogeneous and parallel computing focus, CPU-based and OS-capable devices are of particular interest.
@@
Note that while we use the term ‘DPU’ as described, the nomenclature is in reality inconsistent and depends upon the vendor and product specialization. For example, Intel’s “Infrastructure Processing Unit” (IPU) (not to be confused with an “Intelligence Processing Unit”, which is a fundamentally different device) can be considered a DPU, despite its different name. Furthermore, the term ‘Data Processing Unit’ itself has become overloaded, as it has also been used to describe other devices, such as Processing-in-Memory (PIM) devices, which are not discussed in this work, and the term “DPU” can refer to Deep-Learning Processing Units; neither of these are SmartNICs. In a rather different paradigm, some companies, such as NVIDIA, would argue that the real DPU is actually the specialized processing unit (different from a CPU or a GPU) often included in a SmartNIC, with features irreplaceable by a standard CPU [3]. In our work, we generally adhere to the network card–oriented description of a DPU provided by ServeTheHome [4], discussed in Section 2.2, and use it as a reasonably general differentiation between DPUs and other SmartNICs.
@@
The development of advanced SmartNICs, or DPUs, was started by large cloud service providers (“hyperscalers”) such as Amazon who realized that around 30% of processor cores were being utilized for infrastructure processing [5]. This motivated cloud vendors to develop advanced SmartNICs to offload infrastructure-related tasks related to networking, security, and storage from host CPUs, so that host CPU cores can focus on users’ application workloads. According to the projection of Dell’Oro Group, the traditional NIC market share will be overtaken by SmartNICs by 2027 [6]. We discuss the motivation behind DPU creation and use in Section 2.1, and their evolution in 2.2.
@@
There are many different reasons for developers and professionals to use SmartNICs, the most obvious being their intended use as network infrastructure support devices. For example, one such use case is to perform network offload, where networking tasks, such as packet processing, are delegated to SmartNICs by host CPUs. Their utility extends beyond basic networking functions and is also capable of encompassing security offloading, which is achieved by accelerating cryptographic operations such as those common in a Secure Sockets Layer (SSL), Transport Layer Security (TLS), and encryption. Similarly, SmartNICs are used to offload storage-related operations, as well as to support virtualization tasks.
@@
However, data center infrastructure is not the sole application domain in which SmartNICs are applied, providing the main motivation for this paper. As the role of SmartNICs has grown to capably include offloading virtually any task, treating DPUs as heterogeneous compute nodes, rather than simply accelerators, this survey focuses on works exploring their use in a variety of application-level tasks, in addition to infrastructure acceleration. Such new offload methods might be described by terms such as in-network computing, edge computing, communication acceleration, or simply heterogeneous parallel computing. In both infrastructure and applications, such accelerations are made possible primarily by two things: 1) acceleration and offload of computations and networking functions, and 2) isolation between host and DPU. We discuss offloading briefly in Section 2.3, how data is moved in 2.4, and isolation in 2.5. Section 2.6 discusses challenges regarding their use, and Section 2.7 presents the variety of DPU offerings by technology vendors.
After providing background information on the above topics, the remainder of the survey dives into the current research. Section 3 provides an overview of various programming tools and frameworks for developing application software targeted to DPUs, Section 5 cites a number of broader studies, benchmarking tools, and works which give a big-picture view, and Section 6 discusses various applications of DPUs, including topics such as big data, HPC, and machine learning. As DPUs continue to evolve, their application areas and benchmarks, and the focal points chosen by the research community, will play a pivotal role in shaping the landscape of high-performance computing and data processing. Finally, we conclude with our broad observations and recommendations regarding the field, and which research areas seem most effective, in Section 7.
@@ 2 Background
For those unfamiliar with SmartNICs, a certain amount of background information will help to understand the research we review. These background sections are intended to be a brief overview; other surveys, such as that by Kfoury et al. [1], cover these topics in much greater depth for the benefit of domain experts.
2.1 The Purpose of a DPU
Approximately 30% of the work done in a data center is in handling network infrastructure tasks (called an “infrastructure tax”), which means that it requires almost half again as many servers to get the intended work done; this can be seen as severely bloating the size and cost of data centers. DPUs are increasingly seen as a solution to this problem, and CPUs as insufficient [7].
Heterogeneous computing using emerging Data Processing Units (DPUs) is driven in particular by the necessity for specialized acceleration, enhanced performance, and energy efficiency. DPUs, which are specifically designed for particular types of tasks, optimize the effectiveness of the system by transferring specialized workloads from the host CPU, providing a more energy-efficient solution for data center infrastructure workloads, which also improves cost-efficiency. However, since DPUs have a general-use processor that is also more energy efficient [2], they are capable of accommodating a wide range of workloads and enhancing the overall utilization of resources. In this way, the adaptability and flexibility of heterogeneous computing with DPUs is intended to address the evolving requirements of modern applications [8].
However, the motivation behind the creation of these smart network devices goes deeper. As the world turns to cloud technologies and flexibly hosted web services, there is an increasing need for software-defined networking (SDN)—systems capable of not only specialized, but customizable networking configurations within data centers and supercomputers [9]—perhaps for both infrastructure and tenant applications alike. DPUs will be part of the future infrastructure that powers services such as augmented reality, smart cities, edge computing, big-data analytics, and more [10].
A ‘smarter’ NIC understandably comes with a higher price point and per-node power usage, and whether or not that cost comes out to be worth it is a matter of balance, requiring the consideration of many factors for a given data center, not least of which is the potential for millions of dollars worth of electricity and capital savings for a large data center [2]. However, the desire for software-reconfigurable optimization within the data center appears to generally be worth the cost, given the directions in which technology is advancing; DPU adoption is expected to keep increasing [11].
@@ 2.2 Evolution of NICs and DPUs
Figure 1: Part of the 2021 ServeTheHome NIC Continuum, a useful classification of existing NIC hardware, and which doubles as a representation of its evolution over time [4].
@@
The ‘Foundational NIC’ shown in Figure 1 represents the fundamental level of a network interface. Nearly all contemporary NICs possess certain rudimentary offloads, such as IPv4, IPv6, and TCP/UDP checksum offloads. Foundational NICs are specifically engineered to facilitate low-cost network ports, eschewing many of the advanced offload features that contribute to increased cost of more sophisticated devices. Foundational NICs retain their significance within the industry, but are insufficient in many cases, particularly at higher data rates where the escalation in data processing necessitates enhanced computational capabilities.
Consequently, the prevailing NICs are predominantly mid-tier ‘Offload NICs’, which generally provide networking speeds of about 100Gbps and above [4]. In order to design NICs with such speeds, it becomes necessary to offload more network support functions, generally in the form of on-board hardware accelerators (ASICs, or application-specific integrated circuits) [4]; otherwise, the host CPU would be quite busy trying to keep up.
‘SmartNICs’, however, go beyond simple offloads. These are NICs that possess programmable pipelines, allowing systems and applications to tweak and configure offloads more flexibly. The primary purpose of a SmartNIC, as being part of a server device, is to reduce the burden imposed on the host CPU [4]. “Smart”, in the name, suggests a certain programmability, or at the very least an increased ability to handle complex tasks—both current, and being adaptable to future needs [14].
A ‘DPU’ differs from this paradigm, however, in that it can hardly be considered part of the host server; it is, in and of itself, a mini network endpoint [4, 15], providing ‘off-path’ capabilities (we will define off-path in Section 2.4). The concept of a DPU encompasses the functionality of a SmartNIC and extends beyond it, possessing not only ASIC offload capabilities and the flexible programmable pipelines of a SmartNIC, but also its own memory, storage, and CPUs or FPGAs. This design enables them not only to perform network tasks, but to participate in data processing tasks—further freeing up the host’s compute resources, all while consuming less power than the host [15]. While ServeTheHome further differentiates DPUs from more “Exotic” devices incorporating large Field-Programmable Gate Arrays (FPGA), we lump these in with DPUs for the purposes of this survey. We note also NVIDIA’s ‘SuperNIC’ product line as an important type of SmartNIC; however, as these devices primarily focus on inter-GPU communication, rather than specializing in providing heterogeneous internal computation themselves, we do not discuss them further in this survey [16].
@@ 2.3 Computational Acceleration and Offload
While NICs and SmartNICs already offload some types of work from the host’s CPUs, DPUs are designed to go further, being able to offload virtually any kind of work, with generally the same goal of freeing up host resources to focus on application-level processing. A DPU, being ‘close to’ the network, boasts better network I/O processing performance compared to the host, making it ideal for accelerating network stack and packet processing, session management, distributed storage and file system client execution, erasure coding, network security layers, or acting as a virtual switch [17, 18, 2]. DPUs have also been leveraged in network-edge scenarios for applications such as network awareness, power-efficient edge-to-cloud continuum, etc. [19, 20]. The goal of such offloads is to free up the host’s resources to focus on applications processing tasks and improve system-wide performance by reducing network and data center overhead [21].
And, while DPU acceleration typically refers to offloading these types of network and infrastructure tasks using on-board hardware accelerators (ASICs) [22, 20], this survey especially points out those works which instead show that one can offload generic computations and data manipulation to a DPU, including work in the applications layer, moving some processing from the host to the DPU’s CPUs. Some notable examples from the papers we review include big data processes, machine learning functions and inference, molecular dynamics, and various other forms of scientific computing. A simple illustration of what offloading to a DPU might look like is provided in Figure 2.
@@ 2.4 Data Path
Table 1: Characteristics of on-path and off-path SmartNICs.
When discussing ‘data paths’ in terms of advanced SmartNICs such as DPUs, the term may be more specific than the typical usage. In general, a data path describes the steps through which data travels as it is moved, analyzed, transformed, and processed within a workflow. In a DPU, however, the data path usually traverses both the host and the DPU, and typically can be described one of two ways: ‘on-path’ or ‘off-path’, as shown in Table 1. Understanding the difference between the two is helpful in understanding the various use-cases of SmartNICs and DPUs. In this work, we generally associate the term DPU with off-path devices.
On-path SmartNICs have CPU cores that are situated on the communication path of the network packets traveling to the host, and possess the ability to manipulate each incoming or outgoing packet rapidly. Too much additional processing assigned to these cores can bottleneck the communication path, because all network traffic must pass through them. Examples of on-path SmartNICs include the LiquidIO II CN2350 and CN2360.
By contrast, off-path SmartNICs (typically DPUs) have cores that are not in the direct line of traffic to the host. Their internal NIC can use customizable network forwarding rules to sort between packets that should be sent to the DPU cores first for preprocessing, and those that should be forwarded directly to the host instead. However, direct control over packets and memory will be limited, and will have overheads; any off-path processing which occurs lengthens the data path, and thus time, for packets to make it to the host [23]. Examples of off-path SmartNICs include the Mellanox BlueField and the Broadcom Stingray [24].
In short, a ‘DPU’ is a class of device with certain advanced capabilities, which is usually an ‘off-path’ device, a term which describes both the physical network path within the device, and a paradigm for how the data flows through it. Off-path processing might include, for example, live network flow analysis, data preprocessing steps, etc. Many research projects take advantage of the flexibility of off-path systems, analyzing the data path and data flow of a given application and experimenting with using a SmartNIC to optimize its efficiency, or to save on data preprocessing steps. Others benchmark the DPUs themselves in different path modes [25].
@@ 2.5 Isolation
First, amid the many services or tenants which a server node often has to support, being able to separate or isolate many of the infrastructure tasks from the tenant or application workload, as well as separating those tenant workloads from each other, has the capacity to improve performance. This is not only because it redistributes compute resources, but also because it reduces interference between tasks; CPU context switching can have a high overhead cost, particularly due to cache pollution, and task segregation can help reduce that cost. This problem only worsens when you consider the myriad of additional ASIC resources over which there might be contention [26].
Secondly, a DPU can be used as an extra security layer. Because it functions as a separate device which can process network packets before forwarding them to the host, it provides an added layer of hardware isolation between the network and the host server, effectively isolating infrastructure and other functions performed on the SmartNIC from those performed on the host, which usually amounts to separating ‘control plane’ operations from ‘data plane’ operations, or infrastructure operations from tenant operations, or network access from user applications, thus further protecting the data center from nefarious tenants. And, with more security operations offloaded to ASICs, one could reasonably say that the more computation one does in silicon, instead of software, the more secure the system will be [10].